OpenStudy (anonymous):

What is the probability of rolling a 6-sided die two times and getting a 3 and then a 2?

OpenStudy (mathstudent55):

The two events are independent of each other. That means that the outcome of the first roll does not influence the outcome of the second roll. That being the case, the probability of rolling a 3 followed by a 2 is the product of the two individual probabilities. p(3 then 2) = p(3) * p(2)

OpenStudy (mathstudent55):

What is the probability if rolling a 3? What is the probability of rolling a 2?

OpenStudy (anonymous):

A. 1/3 B. 1/36 C. 1/6 D. 1/18

OpenStudy (mathstudent55):

I see the choices. Can you answer my questions above?

OpenStudy (mathstudent55):

A die has 6 possible outcomes. All 6 outcomes are equally likely. What is the probability of rolling one specific outcome out of the 6?

OpenStudy (anonymous):

i wouldnt know sir/mam, that why im asking you guys for the answer

OpenStudy (mathstudent55):

The idea with OS is to get help, I agree, but it is also to make you understand the material and not to just be told answers. We need to think this together and figure it out.

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

so 1/6

OpenStudy (anonymous):

would be it

OpenStudy (mathstudent55):

Let me give an example with a simpler case. You have a coin. The coin can fall with the heads side up or with the tails side up. There are only two possible outcomes with the flipping of a coin. If I flip the coin, what is the probability it will be a heads?

OpenStudy (anonymous):

1/2 for the coin

OpenStudy (mathstudent55):

In this case, the probability would be 1/2. You are correct. In your case it is 1/6.

OpenStudy (anonymous):

thanks kind person

OpenStudy (mathstudent55):

Right. For the die, any one specific outcome is 1/6.

OpenStudy (mathstudent55):

This is not the final answer.

OpenStudy (anonymous):

really

OpenStudy (anonymous):

so 1/3 :)

OpenStudy (mathstudent55):

No. Read below. Now go back to where I wrote this: p(3 then 2) = p(3) * p(2) p(3) = 1/6 p(2) = 1/6 The probability of rolling a 3 followed by a 2 is the probability of rolling a 3 ___multiplied___ by the probability of rolling a 2. p(3 then 2) = p(3) * p(2) = 1/6 * 1/6

OpenStudy (mathstudent55):

Don't add the individual probabilities. Multiply them instead.

OpenStudy (anonymous):

ohhhhhh its 1/36

OpenStudy (mathstudent55):

You got it. Remember for future problems. If each event is independent of each other, then the individual probabilities multiply.

OpenStudy (anonymous):

sweet

OpenStudy (mathstudent55):

Another way of thinking of this is to find all the possible outcomes of rolling a die twice and then seeing how many are the desired outcome compared to the total number of possible outcomes.

OpenStudy (anonymous):

hmmm ok

OpenStudy (mathstudent55):

If you roll a die twice these are all the possible outcomes: 11, 12, 13, 14, 15, 16 21, 22, 23, 24, 25, 26 31, 32, 33, 34 ,35, 36 41, 42, 43, 44, 45, 46 51, 52, 53, 54, 55, 56 61, 62, 63, 64, 65, 66 As you can see, there are 36 possible different outcomes. You are interested in only one outcome, 32. 1 desired outcome out of 36 possible outcomes means a probability of 1/36

OpenStudy (anonymous):

haha ok thats awesome. you are very helpful

OpenStudy (mathstudent55):

You're welcome. Thanks.